# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v2

@@ Abstract
Tchetgen Tchetgen and Shpitser (2012) introduced an efficient, debiased, and robust influence function-based estimator for the mediation functional, which is the key component in mediation analysis. This estimator relies on the treatment, mediator, and outcome mean mechanisms. However, treating these three mechanisms as nuisance functions and fitting them as accurately as possible may not be the most effective approach. Instead, it is essential to identify the specific functionals and aspects of these mechanisms that impact the estimation of the mediation functional. In this work, we propose a two-stage estimation strategy for certain nuisance functions in the influence function of the mediation functional that are based on these three mechanisms. This strategy is guided by the role those nuisance functions play in the bias structure of the influence function-based estimator for the mediation functional. In the first stage, we estimate two primary nuisance functions, namely the inverse treatment mechanism and the outcome mean mechanism. In the second stage, we leverage these primary functions to estimate two additional nuisance functions that encapsulate the needed information about the mediator mechanism. We propose a nonparametric weighted balancing estimation approach to design the estimator for one of the nuisance functions in Stage 1 and one of in Stage 2, where the weights are designed directly based on the bias of the final estimator for the mediation functional. The remaining two nuisance functions are estimated using standard parametric or nonparametric regression methods. Once all four nuisance functions are obtained, they are incorporated into the influence function-based estimator. We provide a robustness analysis of the proposed method and establish sufficient conditions for consistency and asymptotic normality of our estimator for the mediation functional.
Keywords: Causal Mediation Analysis; Influence Function-Based Estimation; Multiple Robustness; Balancing Estimators.
@@ 1 Introduction
In contrast to simply assessing the overall treatment effect of a treatment variable on an outcome variable of interest, causal mediation analysis delves deeper by posing a “how” inquiry, investigating the underlying mechanisms through which the treatment influences the outcome. Due to this conceptually invaluable focus, it has gained significant popularity in many fields of science including social science, economics, and epidemiology (VanderWeele,, 2015; MacKinnon,, 2012; Hayes,, 2017; Richiardi et al.,, 2013; Ten Have and Joffe,, 2012).
@@
The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model for the data generating process, the direct and indirect effects can be identified by a functional known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those assumptions hold in the setting under study and the parameter is identified, the next question is how to estimate the mediation functional.
@@
There has been several estimation strategies proposed in the literature based on regression and weighting techniques. In their seminal paper, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric efficient and has the desired property that under certain regularity conditions, it remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism (i.e., the conditional distribution of the treatment given the covariates), one describing the mediator mechanism (i.e., the conditional distribution of the mediator variable given covariates and the treatment), and one describing the outcome mean mechanism (i.e., the conditional expectation of the outcome variable given covariates, the treatment, and the mediator), and the model for two out of three should be correctly specified. There are two challenges with their proposed estimator: (1) among the aforementioned nuisance functions, the mediator mechanism is usually a difficult object to estimate, as it may be the density of a multivariate continuous variable conditioned on another multivariate continuous variable; (2) in their estimator, the estimator of the treatment and mediator mechanisms appear in the denominator, and hence can significantly inflate the variance of the estimator, especially when these quantities are small. Therefore, trying to naively fit these conditional distributions as good as possible will not necessarily lead to an estimator with reasonable mean squared error. Instead, it is important to understand what aspects of these objects should be focused on to obtain a better estimate of the mediation functional.
@@
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis, which avoids direct estimation of the treatment and mediator mechanisms. In the first stage, we estimate two basic nuisance functions, namely the inverse of the treatment mechanism and the outcome mean mechanism. In the second stage, using the basic nuisance functions, we estimate two additional nuisance functions that contain the needed information regarding the mediator mechanism. We propose a non-parametric version of the weighted balancing approach of Imai and Ratkovic, (2014) to design the estimator of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide an implementation of our non-parametric estimators with RKHSes as the function classes, and provide closed-form solutions for the nuisance parameters. The other two nuisance functions are obtained using standard parametric or nonparametric regressions. After obtaining all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We provide a robustness analysis of the resulting estimator and demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain a asymptotically unbiased estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest.
The rest of the paper is organized as follows. In Section 2, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. In Section 3, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. We propose our two-stage nuisance function estimation approach as well as its implementation using RKHSes in Section 4. We provide the robustness and asymptotic analyses of the proposed estimator in Section 5. We conclude in Section 6.
@@ 2 Model Description
We consider a model with binary treatment variable $A\in\{0,1\}$, pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. Let $P_{0}$ represent the true underlying distribution from which the independent and identically distributed observations $\{O_{i}:=(X,A,M,Y)_{i}\}_{i=1}^{n}$ are collected. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$, and $Y^{(m)}=Y^{(A,m)}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $p(m\mid a,x)>0$, and $p(a\mid x)>0$.
@@
Intuitively, Part $(i)$ of the assumption requires that the treatment-outcome and mediator-outcome relations are not confounded by unobserved variables, Part $(ii)$ of the assumption requires that the treatment-mediator relation is not confounded by unobserved variables (see Figure 1), and Part $(iii)$ of the assumption, which is a so-called cross-world independence assumption, requires that the potential outcomes of $Y$ and $M$ even for different values of the treatment are conditionally independent. All three parts are untestable and should be decided based on expert’s knowledge and the researcher’s belief about how the variables in the system are generated.
Consider the parameter average treatment effect (ATE) of the treatment $A$ on the outcome $Y$, represented by $\theta_{ATE}=\mathbb{E}[Y^{(1)}-Y^{(0)}]$. The main goal in causal mediation analysis is to quantify the part of the ATE that is mediated by the variable $M$, and the part that is directly reaching the outcome variable. To this end, the ATE can be partitioned as follows (Robins and Greenland,, 1992; Pearl,, 2001).
$\theta_{ATE}$ $=\mathbb{E}[Y^{(1)}-Y^{(0)}]$
$=\mathbb{E}[Y^{(1,M^{(1)})}-Y^{(0,M^{(0)})}]$
$=\underbrace{\mathbb{E}[Y^{(1,M^{(1)})}-Y^{(1,M^{(0)})}]}_{\text{NIE}}+\underbrace{\mathbb{E}[Y^{(1,M^{(0)})}-Y^{(0,M^{(0)})}]}_{\text{NDE}}.$
@@
We notice that by consistency, positivity, and sequential exchangeability, we have
$\mathbb{E}[Y^{(1,M^{(1)})}]=\mathbb{E}[Y^{(1)}]=\mathbb{E}[\mathbb{E}[Y^{(1)}\mid X]]=\mathbb{E}[\mathbb{E}[Y^{(1)}\mid A=1,X]]=\mathbb{E}[\mathbb{E}[Y\mid A=1,X]].$
That is $\mathbb{E}[Y^{(1,M^{(1)})}]$ is identified. Similar argument holds for $\mathbb{E}[Y^{(0,M^{(0)})}]$. The parameter $\theta_{0}=\mathbb{E}[Y^{(1,M^{(0)})}]$ requires more caution. It is known that under the assumptions of our model, this parameter can be identified as follows (Pearl,, 2001; Imai et al.,, 2010).
$\theta_{0}$ $=\mathbb{E}[Y^{(1,M^{(0)})}]$
[…]
@@
The last expression is called the mediation functional (also known as mediation formula). Note that this expression is purely a functional of the observational data. We denote the mediation functional by $\psi_{0}$. Note that under the identification assumptions in our model, $\psi_{0}=\theta_{0}$, but in general, these two parameters are not necessarily equal. In the sequel, we refer to $\psi_{0}$ as the parameter of interest.
@@ 3 Influence Function-Based Estimator
Tchetgen Tchetgen and Shpitser, (2012) showed that in a nonparametric model, the efficient influence function of the mediation functional is given by
$IF_{\psi_{0}}=$ $\frac{I(A=1)p(M\mid A=0,X)}{p(A=1\mid X)p(M\mid A=1,X)}\{Y-\mathbb{E}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{p(A=0\mid X)}\{\mathbb{E}[Y\mid A=1,M,X]-\eta(1,0,X)\}+\eta(1,0,X)-\psi_{0},$
@@
  • $p(A=1\mid X)$, $\forall X\in\mathcal{X}$,
  • $\mathbb{E}[Y\mid A=1,M,X]$, $\forall M\in\mathcal{M},X\in\mathcal{X}$,
  • $p(M\mid A=a,X)$, $\forall M\in\mathcal{M},X\in\mathcal{X},a\in\{0,1\}$.
@@
$\hat{\psi}^{TTS}=$ $\mathbb{E}_{n}\bigg{[}\frac{I(A=1)\hat{p}(M\mid A=0,X)}{\hat{p}(A=1\mid X)\hat{p}(M\mid A=1,X)}\{Y-\hat{\mathbb{E}}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{1-\hat{p}(A=1\mid X)}\{\hat{\mathbb{E}}[Y\mid A=1,M,X]-\hat{\eta}(1,0,X)\}+\hat{\eta}(1,0,X)\bigg{]},$
where $\mathbb{E}_{n}$ is the empirical expectation operator (treating nuisance functions as non-random), and
[…]
The authors showed that this estimator is multiply robust in the following sense.
@@
$\hat{\psi}^{TTS}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  • $\{\hat{p}(A=1\mid X),\hat{\mathbb{E}}[Y\mid A=1,M,X]\}$,
  • $\{\hat{p}(A=1\mid X),\hat{p}(M\mid A=a,X)\}$,
  • $\{\hat{\mathbb{E}}[Y\mid A=1,M,X],\hat{p}(M\mid A=a,X)\}$.
As mentioned in the introduction, there are two concerns of complexity of the nuisance functions and potential large variance about $\hat{\psi}^{TTS}$. In the next section, we propose an alternative estimation strategy that mitigates these issues.

### v3

@@ Abstract
Tchetgen Tchetgen and Shpitser (2012) introduced an efficient, debiased, and robust influence function-based estimator for the mediation functional, which is the key component in mediation analysis. This estimator relies on the treatment, mediator, and outcome mean mechanisms. However, treating these three mechanisms as nuisance functions and fitting them as accurately as possible may not be the most effective approach. Instead, it is essential to identify the specific functionals and aspects of these mechanisms that impact the estimation of the mediation functional. In this work, we consider a specific reparametrization of the likelihood function that requires four nuisance functions. To estimate them, we propose a two-stage estimation strategy guided by the role of the nuisance functions in the bias structure of the influence function-based estimator. In particular, two of the functions are estimated using a novel nonparametric weighted balancing approach that directly targets the bias of the final mediation functional estimator. We show that the resulting estimator is consistent and asymptotically normal under certain conditions and attains multiple robustness against misspecifications of the nuisance functions. In simulations, our estimator demonstrated better stability and reduced bias and mean squared error compared to the original influence function-based estimator and a naive estimator. In an application to NHANES 2013-2014 data, our approach suggests that obesity results in a 60% higher odds of coronary heart disease, with roughly 35% higher odds attributable to the direct pathway independent of Glycohemoglobin.
Keywords: Balancing Estimators; Causal Mediation Analysis; Influence Function-Based Estimation; Multiple Robustness.
@@ 1 Introduction
In contrast to merely assessing the overall effect of a treatment variable on an outcome variable of interest, causal mediation analysis delves deeper by investigating the underlying mechanisms through which the treatment influences the outcome. This unique focus has made it an indispensable tool across diverse fields, including social science, economics, and epidemiology (VanderWeele,, 2015; MacKinnon,, 2012; Hayes,, 2017; Richiardi et al.,, 2013; Ten Have and Joffe,, 2012). The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model, the direct and indirect effects can be identified using a functional of the observed distribution known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those conditions are met in the setting under study and the parameter is identified, the next question is how to estimate the mediation functional.
@@
There has been several estimation strategies proposed in the literature based on regression and weighting techniques (Imai et al.,, 2010; Lange et al.,, 2012; Valeri and VanderWeele,, 2013; VanderWeele et al.,, 2014; Díaz and Hejazi,, 2020; Hejazi et al.,, 2023; Rudolph et al.,, 2023). In their seminal work, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric efficient and has the desired property that, under certain regularity conditions, remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism (i.e., the conditional distribution of the treatment given the covariates), one describing the mediator mechanism (i.e., the conditional distribution of the mediator variable given covariates and the treatment), and one describing the outcome mean mechanism (i.e., the conditional expectation of the outcome variable given covariates, the treatment, and the mediator), and the model for two out of three should be correctly specified. There are three challenges with their proposed estimator: (1) among the aforementioned nuisance functions, the mediator mechanism is usually a difficult object to estimate, as it may be the density of a multivariate continuous variable conditioned on another multivariate continuous variable; (2) in their estimator of the mediation functional, the estimators of the treatment and mediator mechanisms appear in the denominator, and hence can significantly inflate the variance of the estimator, especially when these quantities are small; (3) their choice of nuisance functions is not explicitly guided by the role of nuisance functions in the bias structure of the estimator of the causal parameter of interest. In a later work, Díaz et al., (2021) considered a reparametrization of the likelihood function to avoid the estimation of the mediation mechanism. [^1] Their framework additionally allows for mediator-outcome confounders affected by treatment. Instead, it relies on an additional nuisance function describing the conditional distribution of the treatment given the covariates and the mediator. While their approach also achieves multiple robustness, it shares the same drawbacks as the original formulation of Tchetgen Tchetgen and Shpitser, (2012)’s estimator: their new nuisance function is difficult to estimate in high-dimensional settings, will appear in denominators leading to large variance, and may even be inconsistent with the treatment mechanism under parametric modeling. Moreover, their approach also fails to consider the bias structure of the estimator. In both methods of Tchetgen Tchetgen and Shpitser, (2012) and Díaz et al., (2021), trying to naively fit the nuisance functions as good as possible will not necessarily lead to an estimator with reasonable performance. Instead, it is important to understand what aspects of the nuisance functions should be focused on to obtain a better estimate of the mediation functional.
@@
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis. Our method can be viewed as a reparametrization of the likelihood function, designed to address the estimation challenges discussed above. The key advantage of this reparametrization is that it avoids direct estimation of the treatment and mediator mechanisms. In the first stage, we estimate two basic nuisance functions, namely the inverse of the treatment mechanism and the outcome mean mechanism. In the second stage, these basic nuisance functions are used to construct two additional nuisance functions that contain the needed information regarding the mediator mechanism. We propose a nonparametric weighted balancing approach to design the estimator of the inverse of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide an implementation of our nonparametric estimators with reproducing kernel Hilbert spaces (RKHS) as the function classes, and provide closed-form solutions for the nuisance functions. The other two nuisance functions are obtained using standard nonparametric regressions. After obtaining estimates of all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain an asymptotically unbiased estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest. We evaluate our proposed approach on both synthetic and real data. In Section 7, we apply our method to estimate the effect of obesity on coronary heart disease mediated by Glycohemoglobin, using data from the 2013-2014 cycle of National Health and Nutrition Examination Survey (National Center for Health Statistics,, 2014).
@@ 2 Model Description
In this section, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. We consider a model with binary treatment variable $A\in\{0,1\}$, pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $p(m\mid a,x)>0$, and $p(a\mid x)>0$.
@@
Intuitively, Part (i) of the assumption requires that the treatment-outcome and mediator-outcome relations are not confounded by unobserved variables, Part (ii) requires that the treatment-mediator relation is not confounded by unobserved variables (see Figure 1), and Part (iii), which is a so-called cross-world independence assumption, requires that the potential outcomes of $Y$ and $M$ even for different assignments of the treatment are conditionally independent. All three parts are untestable and should be decided based on expert’s knowledge and the researcher’s belief about how the variables in the system are generated.
Consider the parameter average treatment effect (ATE) of the treatment $A$ on the outcome $Y$, represented by $\theta_{ATE}=E[Y^{(1)}-Y^{(0)}]$. The main goal in causal mediation analysis is to quantify the part of the ATE that is mediated by the variable $M$, and the part that is directly reaching the outcome variable. To this end, the ATE can be partitioned as follows (Robins and Greenland,, 1992; Pearl,, 2001),
$\theta_{ATE}$ $=E[Y^{(1)}-Y^{(0)}]$
$=E[Y^{(1,M^{(1)})}-Y^{(0,M^{(0)})}]$
$=\underbrace{E[Y^{(1,M^{(1)})}-Y^{(1,M^{(0)})}]}_{\text{NIE}}+\underbrace{E[Y^{(1,M^{(0)})}-Y^{(0,M^{(0)})}]}_{\text{NDE}}.$
@@
We notice that under consistency, positivity, and sequential exchangeability, we have
$E[Y^{(1,M^{(1)})}]=E[Y^{(1)}]=E\{E[Y^{(1)}\mid X]\}=E\{E[Y^{(1)}\mid A=1,X]\}=E\{E[Y\mid A=1,X]\}.$
That is $E[Y^{(1,M^{(1)})}]$ is identified. Similar argument holds for $E[Y^{(0,M^{(0)})}]$. The parameter $\theta_{0}=E[Y^{(1,M^{(0)})}]$ requires more caution. It is known that under the assumptions of our model, this parameter can be identified as follows (Pearl,, 2001; Imai et al.,, 2010).
$\theta_{0}$ $=E[Y^{(1,M^{(0)})}]$
$=\int yp(Y^{(1,m)}=y\mid x)p(M^{(0)}=m\mid x)p(x)dydmdx$
$=\int yp(y\mid a=1,m,x)p(m\mid a=0,x)p(x)dydmdx$
@@
The functional $\psi_{0}$ is called the mediation functional (also known as mediation formula). This quantity is purely a functional of the observed data distribution, $P_{0}$. Note that under the identification assumptions in our model, $\psi_{0}=\theta_{0}$, but in general, these two parameters are not necessarily equal. In the sequel, we refer to $\psi_{0}$ as the parameter of interest. Our goal is to estimate this parameter of interest from the independent and identically distributed observations $\{O_{i}:=(X,A,M,Y)_{i}\}_{i=1}^{n}$ drawn from $P_{0}$.
@@ 3 Influence Function-Based Estimator
In this section, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. The authors showed that in a nonparametric model, the efficient influence function of the mediation functional is given by
$IF_{\psi_{0}}=$ $\frac{I(A=1)p(M\mid A=0,X)}{p(A=1\mid X)p(M\mid A=1,X)}\{Y-E[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{p(A=0\mid X)}\{E[Y\mid A=1,M,X]-\eta(1,0,X)\}+\eta(1,0,X)-\psi_{0},$
@@
  1. (i)
    $p(A=1\mid X=\cdot)$,
  2. (ii)
    $E[Y\mid A=1,M=\cdot,X=\cdot]$,
  3. (iii)
    $p(M=\cdot\mid A=a,X=\cdot)$, for $a\in\{0,1\}$.
@@
$\hat{\psi}^{\text{TTS}}=$ $E_{n}\bigg[\frac{I(A=1)\hat{p}(M\mid A=0,X)}{\hat{p}(A=1\mid X)\hat{p}(M\mid A=1,X)}\{Y-\hat{E}[Y\mid A=1,M,X]\}$
$+\frac{I(A=0)}{1-\hat{p}(A=1\mid X)}\{\hat{E}[Y\mid A=1,M,X]-\hat{\eta}(1,0,X)\}+\hat{\eta}(1,0,X)\bigg],$
where $E_{n}$ is the empirical expectation operator (treating nuisance functions as non-random), and
$\hat{\eta}(1,0,X)=\int\hat{E}[Y\mid A=1,M=m,X]\hat{p}(M=m\mid A=0,X)dm.$
The authors showed that this estimator is multiply robust in the following sense,
@@
$\hat{\psi}^{\text{TTS}}$ is unbiased if at least one of the following pairs of nuisance function estimators is correctly specified.
  1. (i)
    $\{\hat{p}(A=1\mid X=\cdot),\hat{E}[Y\mid A=1,M=\cdot,X=\cdot]\}$,
  2. (ii)
    $\{\hat{p}(A=1\mid X=\cdot),\hat{p}(M=\cdot\mid A=a,X=\cdot)\}$,
  3. (iii)
    $\{\hat{E}[Y\mid A=1,M=\cdot,X=\cdot],\hat{p}(M=\cdot\mid A=a,X=\cdot)\}$.
As mentioned in the introduction, the main concerns with $\hat{\psi}^{\text{TTS}}$ are the complexity of the nuisance functions and variance inflation from treatment and mediator mechanisms in the denominator. In the next section, we propose an alternative estimation strategy that mitigates these issues.